A two-digit number is 6 times the units digit. Find the number if the sum of its digit is 6.
Answer by ikleyn(52787)

.
Answer.  24.

Solution

If the decimal form of number is  "ab",  them


10a + b = 6b,   according to the condition.


It implies  10a  = 5b,   or   b = 2a.


The sum of the digit  is 6:

a + b = 6 = a + 2a = 3a  ====>  a = 2;  b = 4.


The number is  "ab" = 24.

Solved.